OpenStudy (anonymous):

verify tat y=x|x| is a solution to y'=2(|y|)^(1/2) >> the absolute value is driving me nuts T.T

OpenStudy (anonymous):

Split it up into piece-wise functions using the definition of absolute value. Then see what happens to each piece.

OpenStudy (anonymous):

Case 1: \(x>0\):\[ |x|x = x^2 \]And then \(y'=2x\), So show \[ 2x=2(|x^2|)^{1/2} \]

OpenStudy (anonymous):

sorry im late. i was forced to eat lunch just now. i tried the first case, but how do i solve the |x| in 2x=2(|x^2|)^1/2 and for the 2nd case, is it -x^2?

hartnn (hartnn):

you need to show that 2x = 2 |x^2|^1/2 so, x^2 is positive always, so |x^2| = x^2, then just use \(\Large (x^m)^n=x^{mn}\) to show that \((x^2)^{1/2}=x\) and yes, in 2nd case, it is -x^2.

OpenStudy (anonymous):

at last!! i got it :D Thanks a LOT i got for 0 for both. so tat means the function is a solution to the DE. but is there a possibility in other question similar to this that i may get one of the cases does not equals to zero? then will the the function be a solution or not?

hartnn (hartnn):

yes, it may happen, and in that case the function will NOT be the solution.